# Bounding User Contributions for User-Level Differentially Private Mean Estimation

### v4

@@ IV Numerical Experiments
In this section, we compare, via numerical experiments, the performance of our clipping strategy for the case when $d=1$, with the widely used strategy in [amin, Sec. 3], which we call the “AKMV” mechanism [^1] The subscript “AKMV” stands for the initials of the last names of the authors of [amin].. We straightforwardly adapt the mechanism in [amin] from the item-level setting to the user-level setting, by considering the contribution of a user to be the sum $\sum_{j\leq m_{\ell}}x_{j}^{(\ell)}$ of the samples it contributes in the user-level setting.
Importantly, the AKMV mechanism uses a privacy budget of $\varepsilon/2$ to first privately estimate the $\left\lceil\left(\frac{2}{\varepsilon}\right)\right\rceil^{\text{th}}$-largest sample, and then uses this value as the clipping threshold. Via entirely heuristic analysis, the authors of [amin] argue the “optimality” of AKMV mechanism. However, importantly, their analysis ignores the error in the private estimation of the the $\left\lceil\left(\frac{2}{\varepsilon}\right)\right\rceil^{\text{th}}$-largest sample, rendering such a claim of optimality incorrect. As we shall see, the clipping strategy in [amin, Sec. 3] performs quite poorly in comparison to our worst-case-error-optimal strategy, in the average case too, when the data samples are drawn i.i.d. from natural distributions.
@@ IV-A Experimental Setup
A natural application of user-level DP mechanisms is to spatio-temporal datasets; similar to previous work [dp_spcom, dp_preprint], we let $U=65$, in line with the largest speed of buses in km/hr, in Indian cities. We consider two collections of numbers of user contributions $\{m_{\ell}:\ \ell\in[L]\}$.
@@ IV-B Performance on Synthetic Datasets
    the “AKMV” mechanism [amin, Sec. 3] that releases $M_{\text{AKMV}}(\mathcal{D})=\widehat{f}_{T}(\mathcal{D})+\mathbf{Z}_{3}$, where $\mathbf{Z}_{3}\sim\text{Lap}\left(\frac{2\Delta_{\widehat{g}_{T}}}{\varepsilon}\right)$, and $\widehat{f}_{T}(\mathcal{D})$ is obtained by clipping the sum of samples of each user to lie in $[0,T]$. Here, $T$ is the estimate of the $\left\lceil\left(\frac{2}{\varepsilon}\right)\right\rceil^{\text{th}}$-largest sample among $\{\sigma^{(\ell)}\}$, which is estimated with privacy budget $\varepsilon/2$.
@@ V Conclusion
In this paper, we revisited the fundamental problem of releasing the sample mean under user-level differential privacy (DP) – a problem that is well-studied in the DP literature, but typically in the presence of additional (e.g., i.i.d.) assumptions on the distribution of data samples. In our work, we make no distributional assumptions on the dataset; instead, we adopt a worst-case approach to studying the estimation error. We then explicitly characterize this error for a broad class of strategies for bounding user contributions by some combination of clipping the numbers of user contributions and/or clipping the sample values themselves. Our analysis allows us to identify the bounding strategy that is optimal with respect to our worst-case error metric. Via numerical experiments, we demonstrate that our strategy is not only optimal for the worst-case error, but also performs much better than the well-known strategy in [amin] for datasets with i.i.d. samples.
@@ Appendix A Characterizing $\ell_{1}$-Projections Onto $\Delta_{\alpha}$
Following Lemma III.1, consider the collection $\mathcal{S}\subseteq\mathbb{R}^{d}$ of points defined as $\mathcal{S}:=\{\mathbf{z}\in\delta_{\alpha}:\ z_{i}\leq a_{i},\ \text{for all $i\in[d]$}\}$. Note that when $a_{i}>\alpha$, for all $i\in[d]$, we have $\mathcal{S}=\delta_{\alpha}$. Observe that $\mathcal{S}$ is a convex subset of $\mathbb{R}^{d}$; hence, by a version of Carathéodory’s Theorem (see the remark after [cover_thomas, Thm. 15.3.5] and [carath]), any point $\mathbf{z}$ in $\mathcal{S}$ can be written as a convex combination of finitely many, in particular, $d$ points in $\mathcal{S}$. Hence, consider any such collection $\mathcal{Z}=\{\mathbf{z}_{1},\ldots,\mathbf{z}_{d}\}$ whose convex hull equals $\mathcal{S}$. The following claim then holds.

### v5

@@ IV Numerical Experiments
In this section, we compare, via numerical experiments, the performance of our clipping strategy for the case when $d=1$, with the widely used strategy in [2, Sec. 3], which we call the “AKMV” mechanism [^1] The subscript “AKMV” stands for the initials of the last names of the authors of [2].. We straightforwardly adapt the mechanism in [2] from the item-level setting to the user-level setting, by considering the contribution of a user to be the sum $\sum_{j\leq m_{\ell}}x_{j}^{(\ell)}$ of the samples it contributes in the user-level setting.
Importantly, the AKMV mechanism uses a privacy budget of $\varepsilon/2$ to first privately estimate the $\left\lceil\left(\frac{2}{\varepsilon}\right)\right\rceil^{\text{th}}$-largest sample, and then uses this value as the clipping threshold. Via entirely heuristic analysis, the authors of [2] argue the “optimality” of AKMV mechanism. However, importantly, their analysis ignores the error in the private estimation of the the $\left\lceil\left(\frac{2}{\varepsilon}\right)\right\rceil^{\text{th}}$-largest sample, rendering such a claim of optimality incorrect. As we shall see, the clipping strategy in [2, Sec. 3] performs quite poorly in comparison to our worst-case-error-optimal strategy, in the average case too, when the data samples are drawn i.i.d. from natural distributions.
@@ IV-A Experimental Setup
A natural application of user-level DP mechanisms is to spatio-temporal datasets; similar to previous work [13, 3], we let $U=65$, in line with the largest speed of buses in km/hr, in Indian cities. We consider two collections of numbers of user contributions $\{m_{\ell}:\ \ell\in[L]\}$.
@@ IV-B Performance on Synthetic Datasets
    the “AKMV” mechanism [2, Sec. 3] that releases $M_{\text{AKMV}}(\mathcal{D})=\widehat{f}_{T}(\mathcal{D})+\mathbf{Z}_{3}$, where $\mathbf{Z}_{3}\sim\text{Lap}\left(\frac{2\Delta_{\widehat{g}_{T}}}{\varepsilon}\right)$, and $\widehat{f}_{T}(\mathcal{D})$ is obtained by clipping the sum of samples of each user to lie in $[0,T]$. Here, $T$ is the estimate of the $\left\lceil\left(\frac{2}{\varepsilon}\right)\right\rceil^{\text{th}}$-largest sample among $\{\sigma^{(\ell)}\}$, which is estimated with privacy budget $\varepsilon/2$.
@@ V Conclusion
In this paper, we revisited the fundamental problem of releasing the sample mean under user-level differential privacy (DP) – a problem that is well-studied in the DP literature, but typically in the presence of additional (e.g., i.i.d.) assumptions on the distribution of data samples. In our work, we make no distributional assumptions on the dataset; instead, we adopt a worst-case approach to studying the estimation error. We then explicitly characterize this error for a broad class of strategies for bounding user contributions by some combination of clipping the numbers of user contributions and/or clipping the sample values themselves. Our analysis allows us to identify the bounding strategy that is optimal with respect to our worst-case error metric. Via numerical experiments, we demonstrate that our strategy is not only optimal for the worst-case error, but also performs much better than the well-known strategy in [2] for datasets with i.i.d. samples.
@@ Appendix A Characterizing $\ell_{1}$-Projections Onto $\Delta_{\alpha}$
Following Lemma III.1, consider the collection $\mathcal{S}\subseteq\mathbb{R}^{d}$ of points defined as $\mathcal{S}:=\{\mathbf{z}\in\delta_{\alpha}:\ z_{i}\leq a_{i},\ \text{for all $i\in[d]$}\}$. Note that when $a_{i}>\alpha$, for all $i\in[d]$, we have $\mathcal{S}=\delta_{\alpha}$. Observe that $\mathcal{S}$ is a convex subset of $\mathbb{R}^{d}$; hence, by a version of Carathéodory’s Theorem (see the remark after [5, Thm. 15.3.5] and [10]), any point $\mathbf{z}$ in $\mathcal{S}$ can be written as a convex combination of finitely many, in particular, $d$ points in $\mathcal{S}$. Hence, consider any such collection $\mathcal{Z}=\{\mathbf{z}_{1},\ldots,\mathbf{z}_{d}\}$ whose convex hull equals $\mathcal{S}$. The following claim then holds.